OpenStudy (kramse18):

The substitution used is u= , formula number . After substitution, ∫sqrte^8x−25dx=∫ ? du.

OpenStudy (freckles):

\[\int\limits (\sqrt{e^{8x}}-25 ) dx \\ \text{ or } \int\limits \sqrt{e^{8x}-25} dx?\]

OpenStudy (kramse18):

the 2nd one

OpenStudy (freckles):

is this fill in the blank or did you mean to put u= something?

OpenStudy (kramse18):

fill in the blank

OpenStudy (freckles):

ok so we do want to integrate this thing then... hmmm... I'm thinking of a trig sub...(not an algebraic one) let's play with this... \[\int\limits \sqrt{(e^{4x})^2-25} dx \\ \sqrt{25} \int\limits \sqrt{ \frac{1}{25} (e^{4x})^2-1} dx \\ 5 \int\limits \sqrt{(\frac{e^{4x}}{5})^2-1} dx \\ \text{ recall } \tan^2(\theta)=\sec^2(\theta)-1 \\ \text{ so \let } \sec(\theta)=\frac{e^{4x}}{5} \\ \]

OpenStudy (freckles):

differentiate both sides

OpenStudy (kramse18):

what did you use as your u?

OpenStudy (freckles):

nothing you need a trig sub to do this one

OpenStudy (freckles):

well you could replace theta with u if you want to

OpenStudy (freckles):

\[\sec(u)=\frac{e^{4x}}{5}\]

OpenStudy (freckles):

still need to differentiate both sides

OpenStudy (freckles):

if you want you can solve that for u

OpenStudy (freckles):

to fill in that space...

OpenStudy (kramse18):

Its asking for a u though

OpenStudy (freckles):

I already replaced theta with u...

OpenStudy (freckles):

and then I just said you could solve that for u if you want

OpenStudy (freckles):

so that you can fill in that blank

OpenStudy (freckles):

I will solve for u if you need... \[\sec(u)=\frac{e^{4x}}{5} \\ u=arcsec(\frac{e^{4x}}{5}) \] could you try differentiate sec(u)=e^(4x)/5

OpenStudy (kramse18):

u=e^(4x)

OpenStudy (freckles):

if you do that as u you will still need to make another substitution later

OpenStudy (freckles):

sec(u)=e^(4x)/5 gets all the substitutions out of the way so the next step is just to integrate

OpenStudy (freckles):

it will be a pretty easy integral by the way

OpenStudy (freckles):

if you are using a table to integrate you have not made that very clear your table is not universal

OpenStudy (kramse18):

yeah, sorry I am having to use tables to integrate. I have the whole integrate. I just needed the fill in the blanks I had posted

OpenStudy (freckles):

your table is not universal though can you post a copy of your table

OpenStudy (freckles):

ok you suggested earlier that we do u=e^(4x) what happens when you find du

OpenStudy (kramse18):

du would be 4e^4xdx

OpenStudy (freckles):

\[\int\limits \sqrt{(e^{4x})^2-5^2} dx \\ \text{ Let } u=e^{4x} \\ du=4 e^{4x} dx \\ \] guess what e^(4x) can be written as u (the one in our du)

OpenStudy (freckles):

\[u=e^{4x} \\ du=4 u dx \\ \frac{1}{4} \frac{1}{u} du=dx\]

OpenStudy (freckles):

so what does our integral look like in terms of u

OpenStudy (freckles):

it should make it clear what formula number you want to use after that

OpenStudy (kramse18):

\[(1/4u)\sqrt{u^2-25}\]

OpenStudy (freckles):

\[\frac{1}{4} \int\limits \frac{\sqrt{u^2-5^2}}{u} du\]

OpenStudy (freckles):

what number does this look like in your book

OpenStudy (freckles):

hint look at the ones between 39 and 42

OpenStudy (freckles):

your formula is in there somewhere

OpenStudy (kramse18):

its 41, thank you

OpenStudy (freckles):

yep replace the u with's e^(4x) and a's with 5 and multiple that whole result by 1/4